# Feel the Difference? A Comparative Analysis of Emotional Arcs in Real and LLM-Generated CBT Sessions

### v1

@@ Abstract
Synthetic therapy dialogues generated by large language models (LLMs) are increasingly used in mental health NLP to simulate counseling scenarios, train models, and supplement limited real-world data. However, it remains unclear whether these synthetic conversations capture the nuanced emotional dynamics of real therapy. In this work, we conduct the first comparative analysis of emotional arcs between real and LLM-generated Cognitive Behavioral Therapy dialogues. We adapt the Utterance Emotion Dynamics framework to analyze fine-grained affective trajectories across valence, arousal, and dominance dimensions. Our analysis spans both full dialogues and individual speaker roles (counselor and client), using real sessions transcribed from public videos and synthetic dialogues from the CACTUS dataset. We find that while synthetic dialogues are fluent and structurally coherent, they diverge from real conversations in key emotional properties: real sessions exhibit greater emotional variability, more emotion-laden language, and more authentic patterns of reactivity and regulation. Moreover, emotional arc similarity between real and synthetic speakers is low, especially for clients. These findings underscore the limitations of current LLM-generated therapy data and highlight the importance of emotional fidelity in mental health applications. We introduce RealCBT, a curated dataset of real CBT sessions, to support future research in this space.
@@
Xiaoyi Wang1, Jiwei Zhang1, Guangtao Zhang2, Honglei Guo3 1Department of Computer Science, Shantou University, Shantou, Guangdong, China 2Department of Automation, Tsinghua University, Beijing, China 3BNRIST, Tsinghua University, Beijing, China
@@ 1 Introduction
Emotion plays a central role in both diagnosis and treatment. In CBT, client emotions reveal cognitive distortions, inform intervention strategies, and signal therapeutic progress. A meaningful session often involves shifts in emotional state—such as the alleviation of distress or the emergence of insight. These emotional trajectories, or emotional arcs, offer a powerful lens through which to assess the depth, authenticity, and therapeutic quality of a dialogue. It is thus important to understand the extent to which synthetic therapy conversations mirror the emotional dynamics observed in real sessions.
@@
To address this gap, we apply emotion trajectory analysis to CBT conversations from two sources: real-world sessions transcribed from public video-sharing platforms (e.g., YouTube and Vimeo) and synthetic dialogues from the CACTUS dataset Lee et al. (2024). We examine emotional arcs from two perspectives: (1) The Emotion Dynamics of Real vs. Synthetic Therapy Dialogues: How does the emotion change in real and LLM-generated CBT dialogues: for entire dialogues, for counselors, and for clients? (2) Real vs. Synthetic Speakers: How do emotional trajectories differ between real and synthetic dialogues when analyzed separately for counselors and clients (i.e., real vs. synthetic counselors; real vs. synthetic clients)?
To answer these questions, we conduct the first in-depth comparison of emotional dynamics between real and LLM-generated CBT dialogues. We adapt the Utterance Emotion Dynamics (UED) framework  Teodorescu and Mohammad (2023); Hipson and Mohammad (2021) to measure emotion trajectories over time within counselor–client interactions.
Our analysis reveals that although synthetic dialogues are structurally fluent, they diverge from real sessions across several key emotional dimensions. Real therapy sessions display greater emotional variability, more emotion-laden language, and more authentic patterns of emotional reactivity and regulation. Additionally, the emotional arc similarity between real and synthetic speakers is low, with correlation scores typically near zero—indicating poor alignment in how emotions evolve throughout the conversation.
@@
Our contributions are twofold: (1) An annotated dataset of real-world CBT dialogues—RealCBT [^1] We are releasing the dataset.; (2) A comparative analysis of emotion dynamics between real and LLM-generated CBT sessions.
@@ 2 Related Work
2.1 Cognitive Behavior Therapy (CBT)
@@ 2.3 Lexicon-based Emotion Dynamics Computation
3 Real-world CBT Dialogue Dataset
@@ 3.1 Data Collection
Due to privacy, ethical, and legal constraints, publicly available CBT dialogue datasets are extremely limited. To enable comparison with LLM-generated data, we collected real-world CBT Dialogues by identifying video clips from public video-sharing platforms such as YouTube and Vimeo. We restricted our search to videos explicitly labeled as CBT-based counseling sessions. For a detailed description of the selection criteria, please see Appendix A.1. In total, we collected 76 video-based CBT dialogues that met these criteria and we name this dataset as RealCBT. Detailed summary statistics of RealCBT is provided in Table 1 in Appendix A.
@@ 3.2 Preprocessing and Transcription
All videos were first converted into a standard MP4 format and then preprocessed to remove non-conversational, user-generated content such as introductory titles, animations, and narration. We used the Transkriptor [^2] https://transkriptor.com/ service to generate initial transcripts for each video. To ensure transcript quality and temporal alignment, each transcript was manually reviewed and corrected to accurately reflect the spoken dialogue in the video. Additionally, we reviewed metadata associated with each video—such as the number of views, likes, and information about the channel owner—to verify that the content originated from credible sources. These indicators collectively suggest that the selected videos represent professionally conducted CBT sessions.
@@ 3.3 Metadata Annotation
Similar to CACTUS, we annotated each dialogue with the following attributes from video metadata: (1) client problem, (2) client gender, and (3) overall client attitude toward the session. We employed three state-of-the-art language models—ChatGPT-4o Mini, Grok-v3, and Gemini 2.0 Flash—to independently infer these tags for each dialogue. A majority voting scheme was then used to determine the final label for each attribute. To evaluate the accuracy of this automated annotation approach, we manually reviewed a randomly selected subset comprising 30% of the dialogues. The automated predictions were fully consistent with the human annotations in this subset, yielding an estimated labeling accuracy of 100%. Detailed distribution of the metadata of RealCBT is provided in Table 1 in Appendix A.
@@ 4 Emotion Dynamics Computation
From the resulting VAD time series, we derive a set of Utterance Emotion Dynamics (UED) metrics Hipson and Mohammad (2021) that capture the shape and character of emotion progression in therapy dialogues: (1) Emotion Mean: The average V/A/D score across a dialogue or speaker’s arc; (2) Emotion Variability: The standard deviation of V/A/D scores, representing emotional fluctuation or richness; (3) Average Displacement Length: Measures the average number of emotion words spoken more or less than usual. Longer displacements suggest stronger or more persistent emotional shifts; (4) Emotion Rise Rate: Captures how quickly and intensely a speaker shifts into an emotional state (emotional reactivity or sensitivity); (5)Emotion Recovery Rate: Indicates how quickly a speaker returns to their baseline emotional state after a shift (emotion regulation).
@@ 5 Experimental Setup
CACTUS consists of 31,564 dialogues with roughly balanced distributions across attributes. In contrast, our real-world CBT dataset is relatively small and exhibits significant imbalance, particularly in terms of client gender and overall attitude toward therapy. Among the 76 collected dialogues, 84.2% of clients are female (64 out of 76), and 90.8% express a positive attitude toward the counseling process (69 out of 76). Only 12 dialogues feature male clients, and just 2 involve clients with a negative attitude—making these subgroups too sparse to support reliable statistical comparisons.
While it is theoretically possible to match gender and attitude proportions when sampling from the synthetic dataset, the severe imbalance in the real dataset would render such comparisons statistically unreliable and potentially misleading.
To ensure a fair and meaningful comparison, we instead adopt a problem-focused sampling strategy. We focus exclusively on the top three most frequent client concerns in our dataset: anxiety and fear (25 dialogues), self-esteem and confidence issues (19), and relationship-related concerns (14). Together, these categories account for 76.3% of the dataset (58 out of 76 dialogues). For a detailed description of the summary statistics and distribution of RealCBT, please see Table A.2 and Table A.2 in Appendix A. We sample synthetic dialogues from CACTUS to match this distribution.
@@ 6 Experimental Results
We analyze and compare the following UED metrics across valence, arousal, and dominance: Emotion Mean, Variability, Displacement Length, Rise and Recovery Rates. To assess statistical significance between real and synthetic groups, we apply the Mann–Whitney U test for non-parametric pairwise comparisons, using a p-value threshold of 0.05. Figure 1 presents boxplots illustrating the distributions of Emotion Mean and Variability for each role—dialogue, counselor, and client—in both real and synthetic sessions. Blue dots indicate the aggregated Emotion Mean and Variability for each affective dimension, summarizing overall emotional trends. Detailed summary statistics for Emotion Mean, Emotion Variability, and all UED metrics across VAD are provided in Table A.2 - A.2 in Appendix A.
@@ 6.1.1 Emotion Mean
As shown in Figure 1, synthetic speakers (the whole dialogue, counselor, and client) generally sound more emotionally elevated or intense than real human speakers. An exception is observed in valence, where real and synthetic clients show similar scores. We use the Mann–Whitney U test to perform pairwise comparisons between real and synthetic groups on Emotion Mean, across 10 independently sampled synthetic datasets. The test results support this observation.
For entire dialogues (as a whole speaker), we found significant differences in arousal in 90.0% of cases (median p-value << 0.001, average effect size: 0.264)
For counselors, we observed significant differences in valence in all comparisons (median p-value: 0.00019, range: $1.361\times 10^{-7}$ to 0.0035, average effect size: 0.357). For clients, we found significant differences in arousal in 70.0% of comparisons (median p-value: 0.02, range: 0.001225 to 0.09596, average effect size: 0.226) and in valence in 90.0% (median p-value: 0.0046, range: $4.178\times 10^{-7}$ to 0.0655), average effect size: 0.260).
@@ 6.1.2 Emotion Variability
For entire CBT dialogues, real sessions exhibited significantly greater variability in arousal than synthetic ones, indicating more fluctuation or richness in real speakers. All 10 Mann–Whitney U comparisons showed significant differences, with a median p-value of $7.677\times 10^{-6}$ (range: $5.053\times 10^{-8}$ to $0.007753$) and an average effect size of 0.836, indicating a strong effect. In contrast, valence and dominance variability were similar between real and synthetic dialogues, with only partial significant differences (i.e., 5/10 in dominance and 7/10 in valence).
For counselors, real speakers also showed significantly greater variability in both arousal and valence compared to synthetic counselors. All comparisons were significant. For arousal, the median p-value was $1.486\times 10^{-5}$ (range: $4.97\times 10^{-7}$ to $0.0004315$), with an average effect size of 0.406. For valence, the median p-value was $5.699\times 10^{-6}$ (range: $1.561\times 10^{-9}$ to $0.001349$), with an average effect size of 0.422. For dominance, real counselors tended to exhibit greater variability, though the differences were not consistently significant across the comparisons.
@@ 6.1.3 Displacement Lengths
We found that speakers in real dialogues talk significantly more emotion words than the ones in synthetic dialogues in both arousal and dominance. In particular, for arousal, all comparisons showed significant differences, with a median p-value of 0.0001618 (range: $4.643\times 10^{-6}$ to 0.001617) and an average effect size of 0.354. For dominance, 70% of comparisons were significant, with a median p-value of 0.02345 (range: 0.0001469 to 0.1449) and an average effect size of 0.232. In contrast, both real and synthetic speakers uttered similar amount of emotion words in dominance with minimal difference—only 1 out of 10 comparisons reaching significance.
Real counselors tend to speak more emotion words than synthetic counselors. 70% of comparisons showed significant differences in both arousal and valence. The median p-values were 0.0233 (arousal) and 0.007435 (valence), with average effect sizes of 0.227 in both cases. In contrast, dominance minimal showed significant differences in only 20% of comparisons, indicating similar amount of emotion words uttered by both real and synthetic counselors.
Real clients utter more emotion words than synthetic ones in the dimensions of arousal and dominance. Arousal showed consistent differences across all comparisons, with a median p-value of 0.0003099 (range: $1.76\times 10^{-6}$ to 0.007527) and an average effect size of 0.335. Dominance showed partial significant differences in 70% of comparisons, with a median p-value of 0.0163 (range: $5.556\times 10^{-5}$ to 0.1787) and an average effect size of 0.239. In contrast, valence showed no significant differences across any comparisons.
@@ 6.1.4 Emotion Rise Rate
We observed that synthetic speakers tend to respond more quickly and intensely to emotional situations than real speakers in dominance (synthetic 0.0332 vs. real 0.0311) and valence (synthetic 0.0255 vs. real 0.0240). In particular, dominance showed relative strong significant differences in 70% of comparisons with the median p-value 0.02603 (range: 0.001542 to 0.2407, average effect size: 0.204). Similarly, valence demonstrate partial significant differences (50% of comparisons, median p-value: 0.04623, effect size: 0.185). However, real speakers seem to be slightly higher arousal scores than synthetic ones with no significant differences.
Real and synthetic counselors have very similar scores in valence (0.0328 vs. 0.0310), arousal (0.0254, 0.0258), and dominance (0.0247, 0.0254), indicating that they react to emotional situations with similar intensity and speed. There were no consistence significant differences observed.
Synthetic clients seem to be more actively respond to emotional situations than real ones in the dimensions of valence (0.0349, 0.0319) and dominance (0.0262, 0.0240). Both dimensions showed significant differences in 60% of comparisons with the median p-value valence 0.02253 (range: 0.004904 to 0.1897, average effect size: 0.197) and dominance 0.02273 (range: 0.001897 to 0.5602, average effect size: 0.191). Both real and synthetic clients reacted similarly in arousal (0.0270, 0.0262) with no significant differences.

### v3

@@ Abstract
Synthetic therapy dialogues generated by large language models (LLMs) are increasingly used in mental health NLP to simulate counseling scenarios, train models, and supplement limited real-world data. However, it remains unclear whether these synthetic conversations capture the nuanced emotional dynamics of real therapy. In this work, we introduce RealCBT, a dataset of authentic cognitive behavioral therapy (CBT) dialogues, and conduct the first comparative analysis of emotional arcs between real and LLM-generated CBT sessions. We adapt the Utterance Emotion Dynamics framework to analyze fine-grained affective trajectories across valence, arousal, and dominance dimensions. Our analysis spans both full dialogues and individual speaker roles (counselor and client), using real sessions from the RealCBT dataset and synthetic dialogues from the CACTUS dataset. We find that while synthetic dialogues are fluent and structurally coherent, they diverge from real conversations in key emotional properties: real sessions exhibit greater emotional variability, more emotion-laden language, and more authentic patterns of reactivity and regulation. Moreover, emotional arc similarity remains low across all pairings, with especially weak alignment between real and synthetic speakers. These findings underscore the limitations of current LLM-generated therapy data and highlight the importance of emotional fidelity in mental health applications. To support future research, our dataset RealCBT is released at https://gitlab.com/xiaoyi.wang/realcbt-dataset.
@@
Xiaoyi Wang1, Jiwei Zhang1, Guangtao Zhang2*, Honglei Guo3* 1Department of Computer Science, Shantou University, Shantou, Guangdong, China 2Department of Automation, Tsinghua University, Beijing, China 3BNRIST, Tsinghua University, Beijing, China
11footnotetext: Corresponding author
@@ 1 Introduction
Emotion plays a central role in both diagnosis and treatment. In CBT, client emotions reveal cognitive distortions, inform intervention strategies, and signal therapeutic progress. A meaningful session often involves shifts in emotional state, such as the alleviation of distress or the emergence of insight. These emotional trajectories, or emotional arcs, offer a powerful lens through which to assess the depth, authenticity, and therapeutic quality of a dialogue. It is thus important to understand the extent to which synthetic therapy conversations mirror the emotional dynamics observed in real sessions.
@@
To address this gap, we curate RealCBT, a dataset of authentic cognitive behavioral therapy dialogues transcribed from public video-sharing platforms (e.g., YouTube and Vimeo). We then conduct a comparative emotion trajectory analysis between these real-world sessions and synthetic dialogues from the CACTUS dataset Lee et al. (2024). We examine emotional arcs from two perspectives: (1) The Emotion Dynamics of Real vs. Synthetic Therapy Dialogues: How does the emotion change in real and LLM-generated CBT dialogues: for entire dialogues, for counselors, and for clients? (2) Emotion Arc Alignment Across Speaker Pairs: To what extent do emotional trajectories align within and across speaker types (specifically among real–real, synthetic–synthetic, and real–synthetic pairings) for both counselors and clients?
To answer these questions, we conduct the first in-depth comparison of emotional dynamics between real and LLM-generated CBT dialogues. We use the well-established Utterance Emotion Dynamics (UED) framework  Teodorescu and Mohammad (2023); Hipson and Mohammad (2021), adapted specifically for CBT dialogues, to measure emotion trajectories over time within counselor–client interactions. Our goal is both to compare real and LLM-generated CBT dialogues and to establish an empirical benchmark for assessing whether synthetic sessions capture the nuanced emotional dynamics of real therapy. This benchmark provides an important foundation for future work exploring more contextualized and neural emotion models.
Our analysis reveals that although synthetic dialogues are structurally fluent, they diverge from real sessions across several key emotional dimensions. Real therapy sessions display greater emotional variability, more emotion-laden language, and more authentic patterns of emotional reactivity and regulation. Emotional arc correlations are low across all pair types, though synthetic–synthetic pairs show slightly higher alignment than real–real or real–synthetic pairs.
@@
Our contributions are threefold:
  • RealCBT dataset: We curate and release RealCBT, one of the first publicly available datasets of authentic CBT dialogues.
  • Empirical benchmark: We establish an empirical benchmark for comparing emotional dynamics in real vs. synthetic therapy dialogues, using the UED framework adapted for CBT.
  • Comparative insights: We identify key divergences between LLM-generated and real CBT sessions, highlighting where current models fail to capture the variability, regulation, and alignment of authentic therapeutic emotion trajectories.
@@ 2 Related Work
2.1 Cognitive Behavior Therapy
@@ 2.3 Lexicon-based Emotion Dynamics Computation
3 RealCBT: Real-world CBT Dialogue Dataset
In this section, we describe the process of curating the RealCBT dataset.
@@ 3.1 Data Collection
Due to privacy, ethical, and legal constraints, publicly available CBT dialogue datasets are extremely limited. To enable comparison with LLM-generated data, we collected real-world CBT Dialogues by identifying video clips from public video-sharing platforms such as YouTube and Vimeo. We restricted our search to videos explicitly labeled as CBT-based counseling sessions. For a detailed description of the selection criteria, please see Appendix A.1. In total, we collected 76 video-based CBT dialogues that met these criteria. Detailed summary statistics of RealCBT is provided in Table 3 in Appendix A.2.
@@ 3.2 Preprocessing and Transcription
All videos were first converted into a standard MP4 format and then preprocessed to remove non-conversational, user-generated content such as introductory titles, animations, and narration. We used the Transkriptor [^1] https://transkriptor.com/ service to generate initial transcripts for each video. To ensure transcript quality and temporal alignment, each transcript was manually reviewed and corrected to accurately reflect the spoken dialogue in the video. Additionally, we reviewed metadata associated with each video (e.g., the number of views, likes, and information about the channel owner) to verify that the content originated from credible sources. These indicators collectively suggest that the selected videos represent professionally conducted CBT sessions.
@@ 3.3 Metadata Annotation
Similar to CACTUS, we annotated each dialogue with the following attributes from video metadata: (1) client problem, (2) client gender, and (3) overall client attitude toward the session. We employed three state-of-the-art language models (i.e., ChatGPT-4o Mini, Grok-v3, and Gemini 2.0 Flash) to independently infer these tags for each dialogue. A majority voting scheme was then used to determine the final label for each attribute. To evaluate the accuracy of this automated annotation approach, we manually reviewed a randomly selected subset comprising 30% of the dialogues. The automated predictions were consistent with the human annotations in this subset, yielding an estimated labeling accuracy of 100%. Detailed distribution of the metadata of RealCBT is provided in Table 4 in Appendix A.2.
@@ 4 Emotion Dynamics Computation
From the resulting VAD time series, we derive a set of UED metrics Hipson and Mohammad (2021) that capture the shape and character of emotion progression in therapy dialogues: (1) Emotion Mean: The average V/A/D score across a dialogue or speaker’s arc; (2) Emotion Variability: The standard deviation of V/A/D scores, representing emotional fluctuation or richness; (3) Average Displacement Length: Measures the average number of emotion words spoken more or less than usual. Longer displacements suggest stronger or more persistent emotional shifts; (4) Emotion Rise Rate: Captures how quickly and intensely a speaker shifts into an emotional state (emotional reactivity or sensitivity); (5)Emotion Recovery Rate: Indicates how quickly a speaker returns to their baseline emotional state after a shift (emotion regulation).
@@ 5 Experimental Setup
We conduct our study using two datasets. CACTUS consists of 31,564 dialogues with broadly balanced distributions across multiple attributes, making it well suited for large-scale statistical analysis. In contrast, RealCBT comprises 76 dialogues collected from real counseling sessions. Although smaller in scale, RealCBT captures authentic therapeutic interactions and naturally reflects demographic and attitudinal patterns found in practice.
To ensure a fair and meaningful comparison between the two datasets, we adopt a problem-focused sampling strategy. Specifically, we focus on the three most frequent client concerns in RealCBT, including anxiety and fear (25 dialogues), self-esteem and confidence issues (19), and relationship-related concerns (14). Together, these account for 76.3% of the dataset (58 out of 76 dialogues). For detailed summary statistics and distributional information, see Table 5 and Table 6 in Appendix A.2. We then sample synthetic dialogues from CACTUS to match this distribution.
While CACTUS is broadly balanced, RealCBT reflects natural demographic and attitudinal skews (e.g., a higher proportion of female clients and generally positive orientations toward therapy). Such contextual characteristics make strict subgroup balancing less stable, further motivating our problem-focused design.
@@ 6 Experimental Results
We analyze and compare the following UED metrics across valence, arousal, and dominance: Emotion Mean, Variability, Displacement Length, Rise and Recovery Rates. Detailed definitions and computation methods for each metric are provided in Section 4. To assess statistical significance between real and synthetic groups, we apply the Mann–Whitney U test ($\alpha=0.05$) to compare real sessions and each of 10 independently sampled synthetic datasets. For robustness, we report median p-values and corresponding effect sizes ($r$). Figure 1 presents boxplots illustrating the distributions of Emotion Mean and Variability for each role (i.e., dialogue, counselor, and client) in both real and synthetic sessions. Blue dots indicate the aggregated Emotion Mean and Variability for each affective dimension, summarizing overall emotional trends. Detailed summary statistics for Emotion Mean, Emotion Variability, and all UED metrics are provided in Table 7 - 10 in Appendix A.3.
@@ 6.1.1 Emotion Mean
As shown in Figure 1, synthetic speakers (whole dialogue, counselor, and client) generally sound more emotionally elevated or intense than real human speakers. An exception is observed in valence, where real and synthetic clients show similar scores. The pairwise comparisons support this observation. For entire dialogues (as a whole speaker), we found significant differences in arousal in 90% of cases (median p-value << 0.001, $r$ = 0.24). For counselors, we observed significant differences in valence in all comparisons (median p-value << 0.001, $r$ = 0.36). For clients, we found significant differences in arousal in 70.0% of comparisons (median p-value << 0.05, $r$ = 0.23) and in valence in 90% (median p-value << 0.05, $r$ = 0.26).
@@ 6.1.2 Emotion Variability
For entire CBT dialogues, real sessions exhibited significantly greater variability in arousal than synthetic ones, indicating more fluctuation or richness in real speakers. The Mann–Whitney U tests showed significant differences, with a median p-value << 0.0001 and $r$ = 0.84, indicating a strong effect. In contrast, valence and dominance variability were similar between real and synthetic dialogues, with only partial significant differences (i.e., 5/10 in dominance and 7/10 in valence).
For counselors, real speakers also showed significantly greater variability in both arousal and valence compared to synthetic counselors. For arousal, all comparisons were significant (median p-value << 0.0001, $r$ = 0.41). For valence, all comparisons were significant (median p-value << 0.00001, $r$ = 0.42). For dominance, real counselors tended to exhibit greater variability, though the differences were not consistently significant across the comparisons.
@@ 6.1.3 Displacement Lengths
We found that speakers in real dialogues talk significantly more emotion words than the ones in synthetic dialogues in both arousal and dominance. In particular, for arousal, all comparisons showed significant differences, with a median p-value << 0.0001 and $r$ = 0.35. For dominance, 70% of comparisons were significant, with a median p-value << 0.05 and $r$ = 0.23. In contrast, both real and synthetic speakers uttered similar amount of emotion words in dominance with minimal difference—only 1 out of 10 comparisons reaching significance.
Real counselors tend to speak more emotion words than synthetic counselors. 70% of comparisons showed significant differences in both arousal and valence. The median p-values were 0.02 (arousal) and 0.007 (valence), with average effect sizes of 0.23 in both cases. In contrast, dominance minimal showed significant differences in only 20% of comparisons, indicating similar amount of emotion words uttered by both real and synthetic counselors.
Real clients utter more emotion words than synthetic ones in the dimensions of arousal and dominance. Arousal showed consistent differences across all comparisons, with a median p-value << 0.001 and $r$ = 0.335. Dominance showed partial significant differences in 70% of comparisons, with a median p-value << 0.05 and $r$ = 0.24. In contrast, valence showed no significant differences across any comparisons.
Arousal Level Real Counselor Snippet Synthetic Counselor Snippet
High "It’s completely understandable to feel overwhelmed! You’ve taken a huge step today—this is a big deal, and you’re doing really well." "Wow! That’s amazing progress! You should be really proud of yourself—this is wonderful!"
Explanation: Demonstrates emotional reappraisal and support, key to emotion regulation. Explanation: General praise, not a deliberate attempt to regulate the client’s emotional state.
Medium "That’s good to hear. Can you walk me through what happened when you started to feel anxious?" "That’s interesting. Let’s explore that feeling more."
Explanation: Shows calibrated empathy by gently probing the client’s state with a tailored follow-up. Explanation: It’s vague and not clearly grounded in what the client just said.
Low "Okay. Let’s keep going." "Thanks for sharing that."
Explanation: Even when emotional intensity is low, the counselor maintains engagement, which contributes to dyadic co-regulation. Explanation: Like a conversational endpoint, with no guidance or co-regulation of emotional flow.
Table 1: Illustrative Counselor Utterances Comparison at Varying Arousal Levels
@@ 6.1.4 Emotion Rise Rate
We observed that synthetic speakers tend to respond more quickly and intensely to emotional situations than real speakers in dominance (synthetic 0.033 vs. real 0.031) and valence (synthetic 0.026 vs. real 0.024). In particular, dominance showed relative strong significant differences in 70% of comparisons with the median p-value << 0.05 and $r$ = 0.2). Similarly, valence demonstrate partial significant differences (50% of comparisons, median p-value << 0.05 and $r$ = 0.19). However, real speakers seem to be slightly higher arousal scores than synthetic ones with no significant differences.
Real and synthetic counselors have very similar scores in valence (0.033 vs. 0.031), arousal (0.025, 0.026), and dominance (0.0247, 0.0254), indicating that they react to emotional situations with similar intensity and speed. There were no consistence significant differences observed.
Synthetic clients seem to be more actively respond to emotional situations than real ones in the dimensions of valence (0.035, 0.032) and dominance (0.026, 0.024). Both dimensions showed significant differences in 60% of comparisons with the median p-value valence << 0.05 and $r$ = 0.197 and dominance << 0.05 and $r$ = 0.19. Both real and synthetic clients reacted similarly in arousal (0.027, 0.026) with no significant differences.